Question


A 33​-term series is written using only the integers +9 and −2. How many such series can be written that have a sum of​0?


Answer


The number of series that can be written is 1 107 568.


Request and Details


Can you please provide a process for solving this question? It may involve permutations and combinations.



1
Expert's answer
2022-05-17T23:30:51-0400

We can form 0, by using 2(+9) and 9(-2).

2x9-9x2=0. This is 11(2 digit "+9" and 9 digit "-2") digits. 33/11=3. So to construct serie whith sum 0, we should distribute 2x3=6 digits "9" by 33 terms. We can do it by "C^6_{33}=\frac{33!}{27!6!}=1107568" ways.